# Enumeration of maps with tight boundaries and the Zhukovsky transformation

### v1

@@ 1.1 Context and motivations
We pursue our investigation, started in [triskell, polytightmaps], of the enumerative properties of maps with tight boundaries. In this episode, we add a new twist to the story, by making an unexpected connection with the theory of topological recursion [Eynard2007, Eynard2016]. In particular, we find a combinatorial interpretation of the Zhukovsky transformation, which was so far used as an analytical tool in this theory.
Let us first recall some context. The enumeration of maps (graphs embedded in surfaces) is a venerable topic in combinatorics, initiated by Tutte in his famous series of “Census” papers, see [Tutte1963] and references therein. Accounts of further developments may be found for instance in [Goulden2004, Schaeffer15]. A fertile connection with random matrices was made in [BIPZ], and it is essentially its ramifications, motivated by the study of 2D quantum gravity—see e.g. [Witten1991, DiFrancesco1995, Ambjoern1997] for overviews—that led to topological recursion.
In the paper [triskell], we found a surprisingly simple formula for the generating function of planar bipartite maps with three tight boundaries (as we will explain in more detail below, a boundary is said tight if it has minimal length in its homotopy class). The tightness property seems to play a role, as without it the corresponding generating function becomes slightly more involved [CoFu12]. One may therefore wonder whether a similar phenomenon occurs for maps of other topologies (higher genus, more boundaries).
This question was already explored in the limit case of tight maps. Colloquially speaking, a tight map is a map in which every face is seen as a boundary, and is forced to be tight. The counting of tight maps was actually investigated first by Norbury [Norbury2010], as tight maps are nothing but fatgraphs describing lattice points in the moduli space of curves. A remarkable quasi-polynomiality phenomenon occurs in this problem: if we denote by $N_{g,n}(b_{1},\ldots,b_{n})$ the number of tight maps of genus $g$ with $n$ boundaries of lengths $b_{1},\ldots,b_{n}$, then Norbury showed that it is a quasi-polynomial of degree $3g-3+n$ in $b_{1}^{2},\ldots,b_{n}^{2}$ depending on the parities, and we gave in [polytightmaps] a bijective construction of these quasi-polynomials in the planar case $g=0$.
In a further paper [NoSc13], Norbury and Scott observed that the quasi-polynomiality phenomenon can be related to a specific feature in topological recursion, namely that the “$x$” meromorphic function attached to the “spectral curve” takes the particular form
@@
This form is nothing but the so-called Zhukovsky transformation [Zhukovskii1910]. The case of tight maps corresponds to taking the second “$y$” function equal to the Zhukovsky variable $z$, but Norbury and Scott showed that the quasi-polynomiality subsists when $y$ is modified.
In this paper, we relate the observation of Norbury and Scott to the enumeration of maps with tight boundaries. At the combinatorial level, the Zhukovsky transformation $z\mapsto x(z)$ corresponds to a substitution: namely we find that it translates the natural idea of transforming a map with tight boundaries into a map with arbitrary boundaries by gluing a “trumpet” onto each boundary (see Figure 2 below for an illustration). By trumpet, we mean a planar map with two boundaries, one of them being tight while the other is arbitrary. We use the term by analogy with [Saad2019], where a similar notion is introduced in the context of hyperbolic geometry. The coefficients $\alpha$ and $\gamma$ appearing in (1) are related to the “basic” generating functions $R$ and $S$ defined in the next section by $\alpha=S$ and $\gamma=R^{1/2}$.
@@
with $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ a quasi-polynomial of degree $3g-3+n$ in $\ell_{1}^{2},\ldots,\ell_{n}^{2}$ depending on the parities, and $\gamma=R^{1/2}$ as before. In particular, in the planar bipartite case, we will derive a fully explicit expression for $T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}$ from the Collet-Fusy formula [CoFu12].
Let us conclude this section by mentioning the recent paper [Budd2023] which introduces the notion of tight boundaries in the context of hyperbolic geometry, and makes the connection with the so-called JT gravity [Saad2019]. A very similar polynomiality phenomenon occurs there, which makes us believe that we are looking at a same reality from different angles.
@@ 1.2 Basic definitions
For $g$ a nonnegative integer, a map of genus $g$ is a cellular embedding of a finite connected multigraph into the closed orientable surface of genus $g$, considered up to isomorphism. By cellular embedding, we mean that the graph (which consists of vertices and edges) is drawn on the surface without edge crossings, and that the connected components of the complement of the graph, called faces, are homeomorphic to open disks. A map of genus zero is said planar.
A map with boundaries is a map in which we mark some faces and vertices and call them boundaries. We shall use the terms boundary-face and boundary-vertex when we want to specify the type of a boundary. A face which is not a boundary is called an inner face. The length of a boundary is by definition equal to its degree (number of incident edge sides) for a boundary-face, and to 00 for a boundary-vertex. A boundary-face is said rooted if one of its incident corners is marked. A boundary-vertex cannot be rooted.
@@
  • The images of the edges pointing away from a given vertex, and listed cyclically in clockwise order, are the edges pointing away from the image vertex, also in cyclic clockwise order. In particular, the graph isomorphism also induces a bijection of the set of faces onto itself, respecting the incidence relations with vertices and edges.
  • The marked elements are preserved.
@@
Figure 1: A toric map, with three distinguished boundary-faces of degree 6, admitting an automorphism group of order $3$. There are only $2$ rooted maps resulting from marking one of the $6$ corners incident to the cyan face, and we must account for this fact by weighing this map by the inverse automorphism group order factor $1/3$.
@@
We think of boundaries as representing holes in the surface. Precisely, we create a puncture inside each boundary-face, while we draw a small circle around each boundary-vertex: the interior of the circle is removed but the edges incident to the vertex remain connected by the circle. See [triskell, Figure 2] for an illustration.
@@
We now introduce two important quantities denoted $R$ and $S$, which are formal power series in $t,t_{1},t_{2},t_{3},\ldots$. The first one $R$ is the generating function of planar maps with two distinguished boundaries of length $1$. The second one $S$ is that of planar maps with two boundaries, one of length $1$ and the other of length 00. It is known [^1] See for instance [census] for a proof in the dual language. Beware that $t$ is a weight per edge in this reference, one has to set $t_{i}=t\,g_{i}$ to match the notations. that, as formal power series, $R$ and $S$ are determined by the following equations:
$\begin{split}R&=t+\sum_{i\geq 1}t_{i}\ [z^{-1}]\left(z+S+\frac{R}{z}\right)^{i-1},\\ S&=\sum_{i\geq 1}t_{i}\ [z^{0}]\left(z+S+\frac{R}{z}\right)^{i-1}.\\ \end{split}$ (3)
@@
In this case, $R$ can also be understood as the generating function of bipartite planar maps with two unrooted boundaries, one of length $2$ and the other of length 00.
@@ 1.3 Outline
The remainder of this paper is organized as follows. Section 2 is devoted to the trumpet decomposition which relates maps with arbitrary boundaries to maps with tight boundaries. We first describe in Section 2.1 the bijection for maps with boundary-faces only, before analyzing its enumerative consequences in Section 2.2. The case of maps having also boundary-vertices is discussed in Section 2.3. In Section 2.4, we derive from the Collet-Fusy formula an explicit expression for the generating function of planar bipartite maps with tight boundaries. Section 3 is devoted to recursion relations satisfied by the generating functions of maps with tight boundaries. These recursion relations consist in adding an extra tight boundary, preserving the genus. We first discuss in Section 3.1 the addition of a boundary-vertex, and then that of a boundary-face in Section 3.2. In Section 3.3, we check that the explicit expression for planar bipartite maps found in Section 2.4 satisfies these recursion relations. Section 4 is devoted to the quasi-polynomiality phenomenon: we first consider maps with boundary-faces only in Section 4.1, revisiting the result of Norbury and Scott. We then treat the case of maps having also boundary-vertices in Section 4.2: the proof is by induction on the number of boundaries, using the recursion relations of Section 3 for the induction step and results from topological recursion for the initialization. For the sake of clarity, we begin with the case of maps with even face degrees. The extension to the general case follows the same strategy but requires extra technical steps which are offloaded to the appendices. Concluding remarks are gathered in Section 5. Additional material is contained in the appendices: Appendix A generalizes the formula of [triskell] for the generating function of planar maps with three tight boundaries (i.e.  “tight pair of pants”) to the non-bipartite case. Appendices B and C contain the proof of the induction step and the initialization step, respectively, for the quasi-polynomiality phenomenon. Appendix D gives a combinatorial expression for the derivatives of inverse functions of several variables, used in Appendix C.
@@ 2 The trumpet decomposition
In this section, we discuss a bijective decomposition of a map with boundaries into a map with tight boundaries and a collection of annular maps with one strictly tight boundary (a trumpet). This generalizes [triskell, Proposition 6.5], which dealt with planar maps with three boundaries, to arbitrary genera and number of boundaries.
@@ 2.1 The decomposition theorem
Let us introduce some notation. A trumpet is a map $M$ of genus 00 with two boundary-faces $f,F$, such that the boundary-face $F$ is rooted, and such that the boundary-face $f$, called the mouthpiece, is unrooted and strictly tight. Recall from Section 1.2 that this means that the contour of $f$ is the unique closed path of minimal length among all closed paths of $M$ that are (freely) homotopic to $f$. By considering a leftmost geodesic from the root corner of the face $F$ aimed towards the face $f$, we may canonically distinguish one corner incident to $f$. We state the following simple but crucial fact.
@@
This lemma is a consequence of the fact that the trumpet $M$ is drawn on a topological cylinder, and is characteristic of this topology: as can be seen for example on Figure 2, tight faces need not have simple boundaries in maps of arbitrary topology (precisely, of negative Euler characteristic $\chi=2-2g-n$, where $g$ is the genus and $n$ is the number of boundaries).
@@
We may assume that the trumpet $M$ is embedded in the plane, with the puncture of the mouthpiece $f$ set at point 00, and the puncture of the rooted face $F$ sent to infinity. We let $\mathbb{D}$ be the unit open disk in $\mathbb{R}^{2}$ and $\mathbb{S}^{1}=\partial\mathbb{D}$ be the unit circle.
The contour of $f$ is a closed path, which we orient arbitrarily and denote by $c$. Plainly, it admits a simple closed subpath $c^{\prime}$ not reduced to a single vertex, and we aim at showing that $c^{\prime}=c$. By the Jordan-Schoenflies theorem, $c^{\prime}$ cuts the plane into two domains $D$ (bounded) and $\mathbb{R}^{2}\setminus\overline{D}$ (unbounded), and moreover, $c^{\prime}$ viewed as a continuous injective mapping $\mathbb{S}^{1}\to\mathrm{Im}(c^{\prime})$ can be extended into a homeomorphism $h:\mathbb{R}^{2}\to\mathbb{R}^{2}$ sending $\mathbb{D}$ to $D$. If 00 does not belong to $D$, it follows $c^{\prime}$ is homotopic to the trivial path in the punctured plane $\mathbb{R}^{2}\setminus\{0\}$, so that $c$ is homotopic to a strictly shorter path, contradicting the tightness assumption of $f$. Therefore, it must be that $0\in D$, so that $c^{\prime}$ is homotopic to $\mathbb{S}^{1}$ and therefore generates the fundamental group $\pi_{1}(\mathbb{R}^{2}\setminus\{0\})\simeq\mathbb{Z}$. Since $c$, being the contour of the face $f$ that contains 00, is also a generator, this implies that $c^{\prime}$ is either homotopic to $c$ or to its reversed path. But the tightness of $f$ shows that the length of $c^{\prime}$ is at least that of $c$, and the only possibility is that $c^{\prime}$ is equal to $c$. ∎
@@
For every $i\in\{1,\ldots,n\}$, we orient the contour of the faces $f^{0}_{i},f_{i}$ in such a way that $f^{0}_{i}$ lies to the right, and $f_{i}$ lies to the left of their oriented contours. Then, since the boundary of $f_{i}$ is a simple closed path, and since $f_{i},f_{i}^{0}$ have same length, one may identify these boundaries, in such a way that the root of $f_{i}^{0}$ and the distinguished corner of $f_{i}$ are matched, by gluing edges sequentially as they appear in contour order, for our choice of orientation. This results in a map $M=\Phi(M_{0},M_{1},\ldots,M_{n})$ with genus $g$ and with the $n$ rooted boundary faces $F_{1},\ldots,F_{n}$ inherited from $M_{1},\ldots,M_{n}$.
@@
Figure 2: An illustration of the decomposition of a map of genus $3$ with three external faces into a map with tight boundaries and with the same topology, and a brassband of three trumpets. We emphasize on this picture that, while the mouthpieces of the trumpets always have simple boundaries by Lemma 2.1, the (tight) external faces of the central map need not have simple boundaries.
This result is illustrated in Figure 2. As already mentioned above, this is a generalization of [triskell, Proposition 6.5]. The main difficulty comes from constructing the reverse mapping $\Phi^{-1}(M)$, where $M$ is a map of genus $g$ with $n$ boundary-faces $F_{1},\ldots,F_{n}$. This consists in considering, for $i\in\{1,2,\ldots,n\}$, the set $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$ of closed paths in $M$ that are homotopic to $F_{i}$, and that have minimal possible length, and to define an order $\prec^{(i)}$ that makes $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$ a lattice. In words, we have $c\prec^{(i)}c^{\prime}$ if $c$ is closer to $F_{i}$ than $c^{\prime}$, although this description requires some interpretation, which is done in [triskell] by working in the universal cover of $M$. Since $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$ is clearly a finite set, this implies that there is a smallest element $c^{i}$ in $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$, that we call the outermost minimal closed path homotopic to $F_{i}$. Cutting along these outermost closed paths, for every $i\in\{1,\ldots,n\}$, produces $n$ trumpets $M_{1},\ldots,M_{n}$ and a remaining map $M_{0}$ with $n$ tight boundaries $f^{0}_{1},\ldots,f^{0}_{n}$, whose lengths match with those of the corresponding mouthpieces.
We will not repeat the detailed argument here and refer the interested reader to [triskell, Section 6.1].

### v2

@@ 1.1 Context and motivations
We pursue our investigation, started in [BGM22, BGM24], of the enumerative properties of maps with tight boundaries. In this episode, we add a new twist to the story, by making an unexpected connection with the theory of topological recursion [EO07, Eyn16]. In particular, we find a combinatorial interpretation of the Zhukovsky transformation, which was so far used as an analytical tool in this theory.
Let us first recall some context. The enumeration of maps (graphs embedded in surfaces) is a venerable topic in combinatorics, initiated by Tutte in his famous series of “Census” papers, see [Tut63] and references therein. Accounts of further developments may be found for instance in [GJ04, Sch15]. A fertile connection with random matrices was made in [BIPZ78], and it is essentially its ramifications, motivated by the study of 2D quantum gravity—see e.g. [Wit91, DGZ95, ADJ97] for overviews—that led to topological recursion.
In the paper [BGM22], we found a surprisingly simple formula for the generating function of planar bipartite maps with three tight boundaries (as we will explain in more detail below, a boundary is said tight if it has minimal length in its homotopy class). The tightness property seems to play a role, as without it the corresponding generating function becomes slightly more involved [CF12]. One may therefore wonder whether a similar phenomenon occurs for maps of other topologies (higher genus, more boundaries).
This question was already explored in the limit case of tight maps. Colloquially speaking, a tight map is a map in which every face is seen as a boundary, and is forced to be tight. The counting of tight maps was actually investigated first by Norbury [Nor10], as tight maps are nothing but fatgraphs describing lattice points in the moduli space of curves. A remarkable quasi-polynomiality phenomenon occurs in this problem: if we denote by $N_{g,n}(b_{1},\ldots,b_{n})$ the number of tight maps of genus $g$ with $n$ boundaries of lengths $b_{1},\ldots,b_{n}$, then Norbury showed that it is a quasi-polynomial of degree 3g3+n3g-3+n in $b_{1}^{2},\ldots,b_{n}^{2}$ depending on the parities, and we gave in [BGM24] a bijective construction of these quasi-polynomials in the planar case $g=0$.
In a further paper [NS13], Norbury and Scott observed that the quasi-polynomiality phenomenon can be related to a specific feature in topological recursion, namely that the “xx” meromorphic function attached to the “spectral curve” takes the particular form
@@
This form is nothing but the so-called Zhukovsky transformation [Zhu10]. The case of tight maps corresponds to taking the second “yy” function equal to the Zhukovsky variable $z$, but Norbury and Scott showed that the quasi-polynomiality subsists when $y$ is modified.
In this paper, we relate the observation of Norbury and Scott to the enumeration of maps with tight boundaries. At the combinatorial level, the Zhukovsky transformation $z\mapsto x(z)$ corresponds to a substitution: namely we find that it translates the natural idea of transforming a map with tight boundaries into a map with arbitrary boundaries by gluing a “trumpet” onto each boundary (see Figure 2 below for an illustration). By a trumpet, we mean a planar map with two boundaries, one of them being tight while the other is arbitrary. We use the term by analogy with [SSS19], where a similar notion is introduced in the context of hyperbolic geometry. The coefficients $\alpha$ and $\gamma$ appearing in (1) are related to the “basic” generating functions $R$ and $S$ defined in the next section by $\alpha=S$ and $\gamma=R^{1/2}$.
@@
with $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ a quasi-polynomial of degree 3g3+n3g-3+n in $\ell_{1}^{2},\ldots,\ell_{n}^{2}$ depending on the parities, and $\gamma=R^{1/2}$ as before. In particular, in the planar bipartite case, we will derive a fully explicit expression for $T^{(0)}_{2\ell_{1},\ldots,2\ell_{n}}$ from the Collet-Fusy formula [CF12].
Let us conclude this section by mentioning the recent paper [BZ23] which introduces the notion of tight boundaries in the context of hyperbolic geometry, and makes the connection with the so-called JT gravity [SSS19]. A very similar polynomiality phenomenon occurs therein, which makes us believe that we are looking at the same reality from different angles.
@@ 1.2 Basic definitions
For $g$ a nonnegative integer, a map of genus $g$ is a cellular embedding of a finite connected multigraph into the closed orientable surface of genus $g$, considered up to orientation-preserving homeomorphism. By cellular embedding, we mean that the graph (which consists of vertices and edges) is drawn on the surface without edge crossings, and that the connected components of the complement of the graph, called faces, are homeomorphic to open disks. A map of genus zero is said planar.
A map with boundaries is a map in which we mark some faces and vertices and call them boundaries. We shall use the terms boundary-face and boundary-vertex when we want to specify the type of a boundary. A face which is not a boundary is called an inner face. The length of a boundary is by definition equal to its degree (number of incident edge sides) for a boundary-face, and to 0 for a boundary-vertex. A boundary-face is said rooted if one of its incident corners is marked. A boundary-vertex cannot be rooted.
@@
  • The images of the edges pointing away from a given vertex, listed cyclically in clockwise order, are the edges pointing away from the image vertex, also in cyclic clockwise order. In particular, the graph isomorphism also induces a bijection of the set of faces onto itself, respecting the incidence relations with vertices and edges.
  • The marked elements, i.e. the (possibly rooted) boundary-faces and boundary-vertices, are preserved.
@@
Figure 1: A toric map, with three distinguished boundary-faces of degree 6, admitting an automorphism group of order 33. There are only 22 rooted maps resulting from marking one of the 66 corners incident to the cyan face, and we must account for this fact by weighing this map by the inverse automorphism group order factor $1/3$.
@@
We think of boundaries as representing holes in the surface. Precisely, we create a puncture inside each boundary-face, while we draw a small circle around each boundary-vertex: the interior of the circle is removed but the edges incident to the vertex remain connected by the circle. See [BGM22, Figure 2] for an illustration.
@@
We now introduce two important quantities denoted by $R$ and $S$, which are formal power series in $t,t_{1},t_{2},t_{3},\ldots$. The first one $R$ is the generating function of planar maps with two distinguished boundaries of length $1$. The second one $S$ is that of planar maps with two boundaries, one of length $1$ and the other of length 0. It is known [^1] See for instance [BDG02] for a proof in the dual language. Beware that $t$ is a weight per edge in this reference, one has to set $t_{i}=t\,g_{i}$ to match the notations. that, as formal power series, $R$ and $S$ are determined by the following equations:
[…]
@@
In this case, $R$ can also be understood as the generating function of bipartite planar maps with two unrooted boundaries, one of length $2$ and the other of length 0.
@@ 1.3 Outline
The remainder of this paper is organized as follows. Section 2 is devoted to the trumpet decomposition which relates maps with arbitrary boundaries to maps with tight boundaries. We first describe in Section 2.1 the bijection for maps with boundary-faces only, before analyzing its enumerative consequences in Section 2.2. The case of maps having also boundary-vertices is discussed in Section 2.3. In Section 2.4, we derive from the Collet-Fusy formula an explicit expression for the generating function of planar bipartite maps with tight boundaries. Section 3 is devoted to recursion relations satisfied by the generating functions of maps with tight boundaries. These recursion relations consist in adding an extra tight boundary, preserving the genus. We first discuss in Section 3.1 the addition of a boundary-vertex, and then that of a boundary-face in Section 3.2. In Section 3.3, we check that the explicit expression for planar bipartite maps found in Section 2.4 satisfies these recursion relations. Section 4 is devoted to the quasi-polynomiality phenomenon: we first consider maps with boundary-faces only in Section 4.1, revisiting the result of Norbury and Scott. We then treat the case of maps having also boundary-vertices in Section 4.2: the proof is by induction on the number of boundaries, using the recursion relations of Section 3 for the induction step and results from topological recursion for the initialization. For the sake of clarity, we begin with the case of maps with even face degrees. The extension to the general case follows the same strategy but requires extra technical steps which are offloaded to the appendices. Concluding remarks are gathered in Section 5. Additional material is contained in the appendices: Appendix A generalizes the formula of [BGM22] for the generating function of planar maps with three tight boundaries (i.e.  “tight pair of pants”) to the non-bipartite case. Appendices B and C contain the proof of the induction step and the initialization step, respectively, for the quasi-polynomiality phenomenon. Appendix D gives a combinatorial expression for the derivatives of inverse functions of several variables, used in Appendix C.
@@ 2 The trumpet decomposition
In this section, we discuss a bijective decomposition of a map with boundaries into a map with tight boundaries and a collection of annular maps with one strictly tight boundary (a trumpet). This generalizes [BGM22, Proposition 6.5], which dealt with planar maps with three boundaries, to arbitrary genera and number of boundaries.
@@ 2.1 The decomposition theorem
Let us introduce some notation. A trumpet is a map $M$ of genus 0 with two boundary-faces $f,F$, such that the boundary-face $F$ is rooted, and such that the boundary-face $f$, called the mouthpiece, is unrooted and strictly tight. Recall from Section 1.2 that this means that the contour of $f$ is the unique closed path of minimal length among all closed paths of $M$ that are (freely) homotopic to $f$. By considering a leftmost geodesic from the root corner of the face $F$ aimed towards the face $f$, we may canonically distinguish one corner incident to $f$. We state the following simple but crucial fact.
@@
This lemma is a consequence of the fact that the trumpet $M$ is drawn on a topological cylinder, and is characteristic of this topology: as can be seen for example in Figure 2, tight faces need not have simple boundaries in maps of arbitrary topology (precisely, of negative Euler characteristic $\chi=2-2g-n$, where $g$ is the genus and $n$ is the number of boundaries).
@@
We may assume that the trumpet $M$ is embedded in the plane, with the puncture of the mouthpiece $f$ set at point 0, and the puncture of the rooted face $F$ sent to infinity. We let $\mathbb{D}$ be the unit open disk in $\mathbb{R}^{2}$ and $\mathbb{S}^{1}=\partial\mathbb{D}$ be the unit circle.
The contour of $f$ is a closed path, which we orient arbitrarily and denote by $c$. Plainly, it admits a simple closed subpath $c^{\prime}$ not reduced to a single vertex, and we aim at showing that $c^{\prime}=c$. By the Jordan-Schoenflies theorem, $c^{\prime}$ cuts the plane into two domains $D$ (bounded) and $\mathbb{R}^{2}\setminus\overline{D}$ (unbounded), and moreover, $c^{\prime}$ viewed as a continuous injective mapping $\mathbb{S}^{1}\to\mathrm{Im}(c^{\prime})$ can be extended into a homeomorphism $h:\mathbb{R}^{2}\to\mathbb{R}^{2}$ sending $\mathbb{D}$ to $D$. If 0 does not belong to $D$, it follows $c^{\prime}$ is homotopic to the trivial path in the punctured plane $\mathbb{R}^{2}\setminus\{0\}$, so that $c$ is homotopic to a strictly shorter path, contradicting the tightness assumption of $f$. Therefore, it must be that $0\in D$, so that $c^{\prime}$ is homotopic to $\mathbb{S}^{1}$ and therefore generates the fundamental group $\pi_{1}(\mathbb{R}^{2}\setminus\{0\})\simeq\mathbb{Z}$. Since $c$, being the contour of the face $f$ that contains 0, is also a generator, this implies that $c^{\prime}$ is either homotopic to $c$ or to its reversed path. But the tightness of $f$ shows that the length of $c^{\prime}$ is at least that of $c$, and the only possibility is that $c^{\prime}$ is equal to $c$. ∎
@@
For every $i\in\{1,\ldots,n\}$, we orient the contour of the faces $f^{0}_{i},f_{i}$ in such a way that $f^{0}_{i}$ lies to the right, and $f_{i}$ lies to the left of their oriented contours. Then, since the boundary of $f_{i}$ is a simple closed path, and since $f_{i},f_{i}^{0}$ have the same length, one may identify these boundaries, in such a way that the root of $f_{i}^{0}$ and the distinguished corner of $f_{i}$ (obtained from the root of $F_{i}$ as explained above) are matched, by gluing edges sequentially as they appear in contour order, for our choice of orientation. This results in a map $M=\Phi(M_{0},M_{1},\ldots,M_{n})$ with genus $g$ and with the $n$ rooted boundary faces $F_{1},\ldots,F_{n}$ inherited from $M_{1},\ldots,M_{n}$.
@@
Figure 2: An illustration of the decomposition of a map of genus $3$ with three external faces into a map with tight boundaries and with the same topology, and a brassband of three trumpets. We emphasize in this picture that, while the mouthpieces of the trumpets always have simple boundaries by Lemma 2.1, the (tight) external faces of the central map need not have simple boundaries.
This result is illustrated in Figure 2. As already mentioned above, this is a generalization of [BGM22, Proposition 6.5]. The main difficulty comes from constructing the reverse mapping $\Phi^{-1}(M)$, where $M$ is a map of genus $g$ with $n$ boundary-faces $F_{1},\ldots,F_{n}$. This consists in considering, for $i\in\{1,2,\ldots,n\}$, the set $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$ of closed paths in $M$ that are homotopic to $F_{i}$, and that have minimal possible length, and to define an order $\prec^{(i)}$ that makes $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$ a lattice. In words, we have $c\prec^{(i)}c^{\prime}$ if $c$ is closer to $F_{i}$ than $c^{\prime}$, although this description requires some interpretation, which is done in [BGM22] by working in the universal cover of $M$. Since $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$ is clearly a finite set, this implies that there is a smallest element $c^{i}$ in $\mathcal{C}_{\mathrm{min}}^{(i)}(M)$, that we call the outermost minimal closed path homotopic to $F_{i}$. Cutting along these outermost closed paths, for every $i\in\{1,\ldots,n\}$, produces $n$ trumpets $M_{1},\ldots,M_{n}$ and a remaining map $M_{0}$ with $n$ tight boundaries $f^{0}_{1},\ldots,f^{0}_{n}$, whose lengths match with those of the corresponding mouthpieces.
We will not repeat the detailed argument here and refer the interested reader to [BGM22, Section 6.1].